OpenStudy (anonymous):

int (4x-2)/(3(x-1)^2)

OpenStudy (anonymous):

I think that's a partial fractions integral. But I will let someone take over here, :-)

OpenStudy (zarkon):

I agree...I would use partial fractions

OpenStudy (zarkon):

you could also do a simple substitution though

OpenStudy (zarkon):

\[u=x-1\] \[du=dx\] \[x=u+1\]

OpenStudy (zarkon):

\[\int \frac{4x-2}{3(x-1)^2}dx\] \[\int \frac{4(u+1)-2}{3(u)^2}du\] expand then integrate term by term